A solution of sulfuric acid (mm = 98.1g/mol) is 21% by weight and has a density of 1.20g/ml. Calculate the molarity and the molality(mol/kg H2O) of sulrfuric acid solution.
Answer #1

21 % by weight means

21 g H2SO4 present in 100 g solution

mass of H2SO4 = 21 g

mass of solvent = 100 - 21 = 79 g

volume of solution = mass / density = 100 / 1.20 = 83.3 mL

now

1) molarity = (W /MW) (1000 / V in mL)

molarity = (21 / 98.1) (1000 / 83.3)

molarity = 2.57 M

2) molality = (W /MW) (1000 / mass of solvent in g)

molality = (21 / 98.1) (1000 / 79)

molality = 2.71 m